A 1.5m-tall cylindrical tank 1.0m in diameter is completely filled with water. how long will it take to empty it with an open gate valve bursting at 300ml/sec?
1
Expert's answer
2019-09-02T09:26:18-0400
"{\rm Rate\: of\: water\: flow}=\frac{\rm{Volume}}{\rm Time}"

"{\rm Time}=\frac{\rm{Volume}}{\rm Rate\: of\: water\: flow}=\frac{\pi R^2h}{v}"

"=\frac{3.14\times (0.5\: m)^2\times 1.5\:m}{300\times 10^{-6}\:{m^3/s}}=3925\:s=65.4\:min"